what concentration ratio of Nh3 to Nh4Cl is necessary to get a buffer solution with a pOH of 4.5? Kb= 1.8x10^-5
Answer #1

PoH = Pkb + log TNHH U] ipko - log (1:8 x16s) INH3 4.74 = 4.5 POH 4.5 = 4.74 + log EN HY Q] Inni -log [NHYU] TNHY = 4.74-4.5